4. (10 points) Solve the given IVP: y + 8y +22y + 20y = 0; y(0) = 0, y(0) = 1, y (0) = 2.
Answer #1

d² y + 8 da The given differential equation is y + 8y +224 +20y = 0 dy +22 dy dx² dx² 20y=0 y= A eme be a trial solution. THence, m² + 6 m+10=0 m=-3ti or -3-i m+2=0 m=-2. Jos m =- 2 - 3 ti, -3-1 Hence, cons which are Thus, general solution of his ydy = (-20) e- dx ( -1) e-2x + (-3x [£3 C2 + (3) Cosa + (-3C3 - (2) sin a #1. Again, differentiating with respect to x, we obtdaz -32 dý = (40) e-2x + [(8C2 -663) Cosx + (8 C₂ +6C2) sina It is that, given y(0) 0 → 4,2-2.0.0-30 [catos O + C, Bin 0 ]--y (O) = 2 D. 2. + e-3.0 - (46) e 2.0 [(862 - 6 C3). Los o + ( 8 C3 + 6. C2). sin o] = ? = 2. - 4 C, + 8 C2 -663 Hence, 3 C,A2 = 1. (-6-2) - - 2 4 2 - 8 43 2 - 3 8 2 1. [-6-8] - 1. [-4-4 - 14 + 8 = -6 Hence, by Cramers rule, Ai Ci 8 2 24 A 42 8 C2=